OpenStudy (anonymous):

I'm so confused: Calculate the product 4/5 x 55 can anyone help?

OpenStudy (gorv):

(4/5)*55=4*11=44

OpenStudy (anonymous):

thx I dint get it at first I got 31

OpenStudy (anonymous):

Answer: = 4/5 × 55/1 = (4 × 55) / (5 × 1) = 220/5 = 44/1 = 44 The answer is 44 Hope this helps! :D